# LLMs Meet Multimodal Generation and Editing: A Survey

### v1

@@ 7.1 Domains
The works that integrate LLMs in the field of audio can be divided into three categories: general audio sounds, music, and speech. Each category poses the challenges and prospects from three pivotal perspectives—generation, understanding, and editing. The summary of the key technical components of LLMs-based approaches for LLMs-related audio tasks is shown in Tab. LABEL:tab:audio_method, and the related audio-language dataset in Tab VIII. The timeline of milestone works is shown in Fig. 8.
@@ 8 Tool-augmented Multimodal Agents
Figure 11: The milestone of multimodal agents that focus on multimodal generation and eidting.
Over the past few months, lots of works [399, 400, 401, 103, 218, 402, 403, 404, 405, 406, 407, 408, 409, 410, 411, 412, 413, 414, 415, 416, 417, 217] known as tool-augmented LLMs have been emerged as a promising direction in human-computer interaction. They empower LLMs to use external tools to enhance the models’ capabilities. Among them, several works [103, 217, 218, 402, 418, 407] which extend LLMs to other modalities beyond the natural language stand out. In contrast to those expert models which only focus on optimizing one specific task, e.g.., image generation or video generation, these pioneering works built upon LLMs can interactively generate or edit images, videos, and audio by invoking corresponding tools. In this section, we mainly focus on reviewing the recent works that aim to extend LLMs to multimodal generation by augmenting them with external tools.
@@ 8.3 Demonstrations
Currently, several multimodal agents [217, 103, 218, 407] can interact with images, either by generating, editing, or understanding them. For example, they augment LLMs with image generation or editing models, such as Stable Diffusion [10], ControlNet [65] and InstructPix2Pix [242], that can create or modify images based on text prompts. As shown in Fig. 12, we make comparisons between Visual ChatGPT, HuggingGPT, InternGPT, and ControlLLM by an example of image generation, which is beneficial to study their capabilities directly. Interestingly, HuggingGPT decomposes the instruction into two tasks and returns two generated images. The generated image from ControlLLM is more aligned to the provided instruction. We can find a slight performance gap between different methods due to the underlying models used in their systems. In contrast to those text2image models like Stable Diffusion [10], multimodal agents are able to generate vivid images in an interactive manner instead of stiffly returning images to the user. In addition, InternGPT and ControlLLM support pointing devices as input to enhance interactivity. As depicted in Fig 13, taking ControlLLM as an example, the user can click the region of interest from the image, and the object lying in this region is then segmented into a mask by using SAM [425]. Next, the user can send an instruction to edit the image such as removing the masked object in the image. Such a manner not only can edit the image more precisely and efficiently, but also improves the rate of success for tool use.
@@ 9 Generative AI Safety
The vulnerability of generative models to attacks or malicious usage presents unique challenges and has attracted significant research attention. Recent research includes optimization-based attacks, prompt-level manipulations, and data poisoning methods: (i) Optimization-based attacks demonstrate the effectiveness of adversarial techniques to degrade model performance[430] or induce biases and harmful outputs[431][432][433][434]. Adversarial attack and detection-based defense research are also conducted in the field of audio and video [435][435][436][437][438][439]. (ii) Prompt-level attacks[440][441][442][443] reveal the risks at the inference level, where human-crafted inputs can bypass safeguards and elicit unsafe outputs, bringing the security challenges. (iii) Data poisoning methods[444][445] expose that models can be manipulated by injecting malicious data inputs when the integrity of training data is broken. These research works underscore the need for comprehensive approaches to enhance model robustness, secure data integrity, and recognize user’s unsafe interactions, addressing the generative AI vulnerabilities.
@@
Following the discussion on techniques to attack large generative models, there are generally two main approaches to defending against undesirable generation content. The first approach involves not modifying the existing parameters of the model but employing detection mechanisms or manipulating the input prompt context. [446] utilizes VLMs to detect and correct hate speech in multimodal memes. [447][448] manipulate LLMs through in-context learning and defending against jailbreaking attacks. For text-to-image generation, an effective and efficient framework named Latent Guard[449] is proposed to detect the unsafe input prompt. Compared to traditional blacklist-based approaches, it is much more robust because the input prompt is checked in a latent space. And the speed is much faster than the LLM-based unsafe detection method. The second strategy enhances safety by aligning the model with human preferences or values using alignment algorithms. The Proximal Policy Optimization (PPO) algorithm, introduced in [450], has been widely used for aligning LLMs. Recently, Direct Preference Optimization(DPO) and related methods [451][452] introduce an improvement over [450] by presenting a more efficient alignment algorithm capable of learning alignment directly from preference data. For image generative models some preference datasets [453] and alignment methods [454] have also been proposed.
@@
Some studies focus on preventing multi-modal generative models from fabricating facts. Powerful generative models that are capable of producing highly realistic videos have attracted significant attention for their potential misuse like Deepfakes [455][456][457]. Deepfake is a technique that can generate realistic content of a certain identity in the form of an image or video. Methods [458][459][460] focus on distinguishing deepfake videos by detecting visual artifacts. While these detection methods are still limited[461] that they do not insist on adversarial attacks so they are hard to work well on new AIGC models that rely on the pattern seen in training.
@@
Generated content can also bring up copyright issues. To solve these problems, we may need to use data attribution and embed watermarks in the generated outputs. Data origin attribution, which involves tracing the model prediction to its original training data, can help find the data causing copyright issues[462][463][464]. Watermark technique [465][466][467] can make generated content distinguishable from real content by integrating the ownership information in the generated content.
@@
In addition, a variety of datasets also have been proposed to evaluate different aspects of generative AI safety. The SafetyBench[468] dataset is a multiple-choice question dataset for assessing unsafe contents which contains 11,435 items spanning 7 safety categories. This effort is complemented by the GOAT-Bench dataset[469] that evaluates unsafe memes with over 6K diverse topics, including implicit hate speech, gender discrimination, and cyberbullying. Furthermore, datasets like ToViLaG and others [470][471][472] have been developed specifically for visual LLMs, shedding light on the challenges of addressing the generation of toxic content, such as offensive texts and inappropriate images. These datasets offer a comprehensive evaluation for further enhancing the safety of generative models, spanning across text and image contents.
@@ 11.2.1 Semantic Audio Synthesis
From text prompts to image sequences, multimodal storytelling can begin with a topic, a script or even a story outline, which serves as the basis for generating complementary modalities such as text, image sequences. For instance, given a prompt about a fantastical adventure, a multimodal generative model could generate vivid imagery depicting characters and scenes, produce an animated video sequence illustrating key events, or compose a thematic musical score to accompany the narrative.

### v2

@@ 7.1 Domains
The works that integrate LLMs in the field of audio can be divided into three categories: general audio sounds, music, and speech. Each category poses the challenges and prospects from three pivotal perspectives—generation, understanding, and editing. The summary of the key technical components of LLMs-based approaches for LLMs-related audio tasks is shown in Table LABEL:tab:audio_method, and the related audio-language dataset in Table VIII. The timeline of milestone works is shown in Fig. 8.
@@ 8 Tool-augmented Multimodal Agents
Figure 11: The milestone of multimodal agents that focus on multimodal generation and editing.
Over the past few months, lots of works [412, 413, 414, 105, 117, 115, 415, 416, 417, 418, 114, 419, 420, 421, 422, 423, 424, 425, 426, 427, 428, 116, 429, 430] known as tool-augmented LLMs have been emerged as a promising direction in human-computer interaction. They empower LLMs to use external tools to enhance the models’ capabilities. Among them, several works [105, 116, 117, 115, 118, 114] which extend LLMs to other modalities beyond the natural language stand out. In contrast to those expert models which only focus on optimizing one specific task, e.g.., image generation or video generation, these pioneering works built upon LLMs can interactively generate or edit images, videos, and audio by invoking corresponding tools. In this section, we mainly focus on reviewing the recent works that aim to extend LLMs to multimodal generation by augmenting them with external tools.
@@ 8.3 Demonstrations
Currently, several multimodal agents [116, 105, 117, 114] can interact with images, either by generating, editing, or understanding them. For example, they augment LLMs with image generation or editing models, such as Stable Diffusion [30], ControlNet [67] and InstructPix2Pix [249], that can create or modify images based on text prompts. As shown in Fig. 12, we make comparisons between Visual ChatGPT, HuggingGPT, InternGPT, and ControlLLM by an example of image generation, which is beneficial to study their capabilities directly. Interestingly, HuggingGPT decomposes the instruction into two tasks and returns two generated images. The generated image from ControlLLM is more aligned to the provided instruction. We can find a slight performance gap between different methods due to the underlying models used in their systems. In contrast to those text2image models like Stable Diffusion [30], multimodal agents are able to generate vivid images in an interactive manner instead of stiffly returning images to the user. In addition, InternGPT and ControlLLM support pointing devices as input to enhance interactivity. As depicted in Fig. 13, taking ControlLLM as an example, the user can click the region of interest from the image, and the object lying in this region is then segmented into a mask by using SAM [437]. Next, the user can send an instruction to edit the image such as removing the masked object in the image. Such a manner not only can edit the image more precisely and efficiently, but also improves the rate of success for tool use.
@@ 9 Generative AI Safety
The vulnerability of generative models to attacks or malicious usage presents unique challenges and has attracted significant research attention. Recent research includes optimization-based attacks, prompt-level manipulations, and data poisoning methods: (i) Optimization-based attacks demonstrate the effectiveness of adversarial techniques to degrade model performance [442] or induce biases and harmful outputs [443, 444, 445, 446]. Adversarial attack and detection-based defense research are also conducted in the field of audio and video  [447, 447, 448, 449, 450, 451]. (ii) Prompt-level attacks [452, 453, 454, 455] reveal the risks at the inference level, where human-crafted inputs can bypass safeguards and elicit unsafe outputs, bringing the security challenges. (iii) Data poisoning methods [456, 457] expose that models can be manipulated by injecting malicious data inputs when the integrity of training data is broken. These research works underscore the need for comprehensive approaches to enhance model robustness, secure data integrity, and recognize user’s unsafe interactions, addressing the generative AI vulnerabilities.
@@
Following the discussion on techniques to attack large generative models, there are generally two main approaches to defending against undesirable generation content. The first approach involves not modifying the existing parameters of the model but employing detection mechanisms or manipulating the input prompt context.  [458] utilizes VLMs to detect and correct hate speech in multimodal memes.  [459, 460] manipulate LLMs through in-context learning and defending against jailbreaking attacks. For text-to-image generation, an effective and efficient framework named Latent Guard [461] is proposed to detect the unsafe input prompt. Compared to traditional blacklist-based approaches, it is much more robust because the input prompt is checked in a latent space. And the speed is much faster than the LLM-based unsafe detection method. The second strategy enhances safety by aligning the model with human preferences or values using alignment algorithms. The Proximal Policy Optimization (PPO) algorithm, introduced in  [462], has been widely used for aligning LLMs. Recently, Direct Preference Optimization(DPO) and related methods  [463, 464] introduce an improvement over  [462] by presenting a more efficient alignment algorithm capable of learning alignment directly from preference data. For image generative models some preference datasets  [465] and alignment methods  [466] have also been proposed.
@@
Some studies focus on preventing multi-modal generative models from fabricating facts. Powerful generative models that are capable of producing highly realistic videos have attracted significant attention for their potential misuse like Deepfakes  [467, 468, 469]. Deepfake is a technique that can generate realistic content of a certain identity in the form of an image or video. Methods  [470, 471, 472] focus on distinguishing deepfake videos by detecting visual artifacts. While these detection methods are still limited [473] that they do not insist on adversarial attacks so they are hard to work well on new AIGC models that rely on the pattern seen in training.
@@
Generated content can also bring up copyright issues. To solve these problems, we may need to use data attribution and embed watermarks in the generated outputs. Data origin attribution, which involves tracing the model prediction to its original training data, can help find the data causing copyright issues [474, 475, 476]. Watermark technique  [477, 478, 479] can make generated content distinguishable from real content by integrating the ownership information in the generated content.
@@
In addition, a variety of datasets also have been proposed to evaluate different aspects of generative AI safety. The SafetyBench [480] dataset is a multiple-choice question dataset for assessing unsafe contents which contains 11,435 items spanning 7 safety categories. This effort is complemented by the GOAT-Bench dataset [481] that evaluates unsafe memes with over 6K diverse topics, including implicit hate speech, gender discrimination, and cyberbullying. Furthermore, datasets like ToViLaG and others  [482, 483, 484] have been developed specifically for visual LLMs, shedding light on the challenges of addressing the generation of toxic content, such as offensive texts and inappropriate images. These datasets offer a comprehensive evaluation for further enhancing the safety of generative models, spanning across text and image contents.
@@ 11.2.1 Semantic Audio Synthesis
From text prompts to image sequences, multimodal storytelling can begin with a topic, a script, or even a story outline, which serves as the basis for generating complementary modalities such as text and image sequences. For instance, given a prompt about a fantastical adventure, a multimodal generative model could generate vivid imagery depicting characters and scenes, produce an animated video sequence illustrating key events, or compose a thematic musical score to accompany the narrative.